BaconLover:

@warriorz13 Sam is observing the velocity of a car at different times. After three hours, the velocity of the car is 53 km/h. After six hours, the velocity of the car is 62 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. Part B: How can you graph the equation obtained in Part A for the first seven hours?

BaconLover:

@Warriorz13

Warriorz13:

To answer the question above, determine the velocity of the car at time zero. To do such, determine the increase in velocity every hour (acceleration). a = (62 - 53) / (6 -3) = 3 km/ h^2 Calculate for the initial velocity (v) by using either of the velocities paired with the corresponding time, v = 53 km/h - (3 km/h^2) x 3 h = 44 km/h The equation is y = 44 + 3x. Where x is the time in hours, y is the velocity in km/h.

BaconLover:

Omg thank you!! I should marry u now!!

Warriorz13:

Maybe ;)

BaconLover:

Oh :)

Warriorz13:

Then I believe that last sentence should explain how to graph it.

Warriorz13:

:) <3

BaconLover:

i aint graphing this *** lol

Warriorz13:

x'D